Hi! Please, I would so appreciate some help with this problem!
The mean salary of 8 employees is $27,850, and the mean salary of 10 others is $31,600. Find the mean salary of the 18 employees. (Round answer to the nearest cent.)
???

Answer by robertb(4012)

Let E = sum of the salaries of the 8 employees, and
T = sum of the other 10 employees' salaries.
==> E/8 = 27,850, and T/10 = 31,600
==> E = 222,800, and T = 316,000
==> sum of salaries of all employees is E + T = 538,800
==> mean salary for all 18 employees is (E + T)/18 = 538,800/3 = 29,933.33.